OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 33 feet and the child ride has a radius of 22 feet, what is the ratio between the volumes of each ride?

OpenStudy (anonymous):

A) 27:8 B) 9:6 C) 9:4 D) 3:2

OpenStudy (anonymous):

hi taz

OpenStudy (anonymous):

isis go do your home work other wise im taking your tv from the livning room

OpenStudy (anonymous):

your answer is a

OpenStudy (anonymous):

i dont believe you

OpenStudy (anonymous):

how is it a?

OpenStudy (anonymous):

thank you^

OpenStudy (agent0smith):

Volume of a cone \[\Large V = \frac{ 1 }{ 3 } \pi r^ 2 h\]If the radii are in proportion 33:22, then the heights (h) must also be in that same proportion (since they're similar cones)

OpenStudy (agent0smith):

33:22 is the same as 3:2 (divide them both by 11). Which means the heights must be in ratio 3:2, so you can use 3h as the height of the 33ft cone, and 2h as the height of the 22 ft cone. So plug them in. r=33 and 3h \[\Large V = \frac{ 1 }{ 3 } \pi \times 33^ 2 \times (3 h)\] r=22 and 2h \[\Large V = \frac{ 1 }{ 3 } \pi \times 22^ 2 \times (2 h)\] Solve simplify those, then divide the first by the second.

OpenStudy (agent0smith):

That should say "simplify those, then divide the first by the second"

OpenStudy (anonymous):

2.25*1.5h

OpenStudy (agent0smith):

I don't know what that is.. show your work in steps first simplify this one \[\Large V = \frac{ 1 }{ 3 } \pi \times 33^ 2 \times (3 h)\]

OpenStudy (anonymous):

i think the answer is 9:4

OpenStudy (agent0smith):

Do the steps i showed above... it doesnt look like 9:4 find this \[\huge \frac{ \frac{ 1 }{ 3 } \pi \times 33^ 2 \times (3 h)}{\frac{ 1 }{ 3 } \pi \times 22^ 2 \times (2 h) }\] notice you can cancel off a lot of common factors.

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

Yes what...?

OpenStudy (anonymous):

sorry didnt read it all the way threw .. 2.25/1.5h

OpenStudy (agent0smith):

That's not it...

OpenStudy (agent0smith):

There should be no h, they should both cancel off. But 2.25/1.5 isn't right.

OpenStudy (agent0smith):

@Taz_kapadia did you manage to finish?